I am suppose to solve the equation for x: x^2 + 6X = -5
This is what I have done:
X^2 + 6X + 5 = 0
X = 6
X = -5?
i AM CONFUSED
Thanks,
lackinmath

Answer by checkley71(8403)

X^2+6X+5=0
(X+5)(X+1)=0
X+5=0
X=-5
X+1=0
X=-1
THEREFORE THE ANSWERS ARE -5 & -1